OpenStudy (anonymous):

if "i" is raised to an odd power, then it cannot simplify to be a)-1 b)-i c)i

OpenStudy (anonymous):

-1

OpenStudy (lgbasallote):

\[i^1 = i\] \[i^2 = -1\] \[i^3 = i^2 \times i = -1 \times i = -i\] \[i^4 = (i^2)^2 = (-1)^2 = 1\] \[i^5 = (i^2)^2 \times i = 1 \times i = i] the process repeats

OpenStudy (lgbasallote):

\[i^5 = (i^2)^2 \times i = 1 \times i = i\]

OpenStudy (jamesj):

@sauravshakya, please do not give answers with no explanation. It is contrary to the ethos of this site and against our code of conduct: http://openstudy.com/code-of-conduct

OpenStudy (klimenkov):

Odd number \(n\) is a number that can be written in form \(n=2k-1, \ k\) is integer. \(i^{2k-1}=i^{2k}\cdot i^{-1}\) \(i^{2k}=(i^2)^k=(-1)^k\) \(i^{-1}=\frac 1 i=\frac {i} {i\cdot i}=-i\) So, \(i^{2k-1}=i^{2k}\cdot i^{-1}=(-1)^k\cdot (-i)=(-1)^{k+1}\cdot i\) For any integer \(k\) it will be a complex number, but not an integer.

OpenStudy (anonymous):

ok. My mistake